# TraCS: Trajectory Collection in Continuous Space under Local Differential Privacy

### v1

@@ 3.5.3. Other Piecewise-based Mechanisms
In Definition 3.1 and Definition 3.3, we use specific instantiations for the parameters $p$ and $[l,r)$ in the piecewise-based mechanism. Besides these instantiations, TraCS can be integrated with any other piecewise-based mechanisms (Wang et al., 2019; Li et al., 2020; Ma et al., 2024; Liu et al., 2024; Zhao et al., 2021) by redesigning them. For example, SW (Li et al., 2020) is a famous mechanism proposed for distribution estimation; it is defined on $[0,1)\to[-b,1+b)$, with $b$ as a parameter. It uses different $p$ and $[l,r)$, and employs a strategy of maximizing the mutual information between the perturbed and private data to determine $b$.  We can redesign SW for the direction perturbation and distance perturbation in TraCS-D and TraCS-C. Appendix B.4 provides the detailed redesign of SW for them and shows intuitive comparisons with Definition 3.1 and Definition 3.3.
3.5.4. Other Shapes of Location Space
@@
TraCS-D for circular area. When the location space is circular with radius $R$, TraCS-D has a more straightforward design. The main insight is that such a location space $\mathcal{S}_{R}$ has a natural representation in polar coordinates, i.e. $\mathcal{S}_{R}=\{(\varphi,r):\varphi\in[0,2\pi),r\in[0,R)\}$. Any location $\tau_{i}\in\mathcal{S}_{R}$ corresponds to a pair of coordinates $(\varphi_{i},r_{i})$, and this representation is unique and reversible.  In this case, the direction space $\mathcal{D}_{\varphi}=[0,2\pi)$ and the distance space $\mathcal{D}_{r}=[0,R)$ are independent, so we can apply $\mathcal{M}_{\circ}$ to $\varphi_{i}$ and $\mathcal{M}_{-}$ to $r_{i}/R$.
4. Evaluation
This section evaluates the performance of TraCS 222 Code, experiments, and slides are available at https://github.com/ZhengYeah/TraCS. on both continuous and discrete spaces.
4.1. Results on Continuous Space
The strawman method, designed for direction perturbation, can be combined with the same distance perturbation in TraCS-D to ensure LDP for continuous space $\mathcal{S}\subset\mathbb{R}^{2}$. We will still refer to this extended method as the strawman method and compare it with TraCS in continuous space. Additionally, we will examine the effects of different piecewise-based mechanisms and evaluate the performance of TraCS in circular areas.
4.1.1. Setup
(a) $\mathcal{S}=[0,1)\times[0,1)$
(b) $\mathcal{S}=[0,2)\times[0,10)$
Figure 4. Comparison of TraCS and the strawman method.
[…]
Comparison of TraCS and the strawman method.
Figure 5. Conditions TraCS-D outperforms TraCS-C.
Figure 6. TraCS-D vs $k$-RR + uniform sampling.
[…]
TraCS-D vs $k$-RR + uniform sampling.
TraCS requires the following parameters: the location space $\mathcal{S}$, the private trajectory $\mathcal{T}$, the privacy parameter $\varepsilon$, and $\varepsilon_{d}$ for direction perturbation in TraCS-D. The strawman method also requires an instance of $k$ in $k$-RR. We use configurations:
  • Location space $\mathcal{S}$: The default location space is $[0,1)^{2}$. We consider a larger space $\mathcal{S}=[0,2)\times[0,10)$ for comparison.
  • $k$-RR: We set $k=6$ as the strawman method’s default value, and consider $3$-RR and $12$-RR for comparison.
Trajectory utility metric. We follow the commonly used Euclidean distance to measure the utility of the perturbed trajectory. The distance between two locations $\tau_{i}$ and $\tau^{\prime}_{i}$ is defined as $\|\tau_{i}-\tau^{\prime}_{i}\|_{2}$. Therefore, given a private trajectory $\mathcal{T}$ and a perturbed trajectory $\mathcal{T}^{\prime}$, the average error (AE) among all locations in the trajectory is:
where $|\mathcal{T}|$ is the number of locations in the trajectory. A smaller $AE$ indicates better utility of the perturbed trajectory.
To eliminate the impact of specific trajectories, we generate $100$ random trajectories for each experiment and report the average $AE$ over all the trajectories.
4.1.2. Error Comparison
Figure 4(b) presents the error comparison between TraCS and the strawman method. We can observe the error advantage of TraCS over the strawman method as the privacy parameter $\varepsilon$ increases. Specifically, in Figure 4(a), TraCS-C exhibits smaller $AE$ when $\varepsilon\gtrsim 3$, and TraCS-D exhibits smaller $AE$ when $\varepsilon\gtrsim 5$. When the location space expands to $[0,2)\times[0,10)$, i.e. Figure 4(b), the advantage of TraCS becomes more significant, outperforming the strawman method at smaller $\varepsilon$ values.  Among TraCS, TraCS-C consistently has a smaller $AE$ than TraCS-D. This is because TraCS-D often deals with larger distance spaces $\mathcal{D}_{r(\varphi)}$, which also magnifies the error of direction perturbation.  Statistically, the mean $AE$ of TraCS-D is $91.1\%$ of the strawman method in Figure 4(a) and $86.6\%$ in Figure 4(b) across all $\varepsilon$ values. The mean $AE$ of TraCS-C is $75.5\%$ of the strawman method in Figure 4(a) and $64.0\%$ in Figure 4(b) across all $\varepsilon$ values.
Conditions TraCS-D outperforms TraCS-C. The error of TraCS-C depends on the size of the distance spaces $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$, while the error of TraCS-D depends only on the size of the distance space $\mathcal{D}_{r(\varphi)}$. Therefore, if $\mathcal{D}_{r(\varphi)}$ is majorly smaller than $\mathcal{D}_{a}$ or $\mathcal{D}_{b}$, TraCS-D will outperform TraCS-C.  We conduct experiments to verify this condition. Specifically, we set $\mathcal{S}=[0,2)\times[0,10)$, and choose a private trajectory along the $x$-axis, i.e. the short side of $\mathcal{S}$. Then we perturb the trajectory $1000$ times using TraCS-C and TraCS-D and calculate the average $AE$. In this case, $|\mathcal{D}_{r(\varphi)}|\approx 2$, much smaller than $|\mathcal{D}_{b}|=10$. Consequently, TraCS-D exhibits smaller $AE$ than TraCS-C across all the $\varepsilon$ values, as shown in Figure 6. In this experiment, the mean $AE$ of TraCS-D is $49.8\%$ of TraCS-C across all the $\varepsilon$ values.
Figure 7. Comparison between different piecewise-based mechanisms.
Figure 8. Example: TraCS-D for circular area.  Private location $\tau=(\pi,0.5)$.
Error of direction perturbation. Figure 6 shows the average error of direction perturbation in TraCS-D compared with the strawman method. We collect the private directions of the trajectories in the location space $\mathcal{S}=[0,1)\times[0,1)$, and perform direction perturbation using TraCS-D and the strawman method. For the strawman method, we use $k$-RR with $k=3,6,12$. This experiment reflects empirical error results of mechanism $\mathcal{M}_{\circ}$ in Definition 3.1 and $k$-RR + uniform sampling.  We can observe an exponentially decreasing error of TraCS-D as $\varepsilon$ increases, as stated in Theorem 3.5. In contrast, the strawman method exhibits an eventually stable error as $\varepsilon$ increases, due to the inherent inner-sector error of $k$-RR.
4.1.3. Effect of Different Piecewise-based Mechanisms
Figure 8 compares the performance of different piecewise-based mechanisms for TraCS-D. In addition to the mechanisms defined in Definition 3.1 and Definition 3.3, we name the redesigned SW for TraCS-D as TraCS-D (R-SW), as detailed in Appendix B.4.  We can observe that TraCS-D consistently exhibit smaller errors than TraCS-D (R-SW) across all the $\varepsilon$ values. Statistically, the mean $AE$ of TraCS-D is $52.6\%$ of TraCS-D (R-SW) across all the $\varepsilon$ values. This difference comes from the MSE of their sampling distributions, where Definition 3.1 and Definition 3.3 have smaller MSE than R-SW.
4.1.4. TraCS-D for Circular Area
Figure 8 illustrates an example of TraCS-D applied to a circular area $\mathcal{S}=[0,2\pi)\otimes[0,1)$. We set the private location $\tau$ as $(\pi,0.5)$ and the privacy parameter $\varepsilon=5$, then collect $50$ random samples of perturbed locations using TraCS-D. In this case, the dominant area is $[0.87\pi,1.13\pi)\otimes[0.32,0.67)$, as determined by the perturbation mechanisms in Definition 3.1 and Definition 3.3. We can observe that the perturbed locations are concentrated in the dominant area.
4.1.5. Choose TraCS-D or TraCS-C
4.2. Results on Discrete Space
4.2.1. Setup
(a) $m=10\times 10$
(b) $m=50\times 50$
Figure 9. Comparison on synthetic datasets. $m$ is the number of discrete points in the location space.
[…]
Comparison of TraCS and the strawman method.
We compare the following methods:
  • ATP (Zhang et al., 2023): We set the $k$-RR parameter $k=6$ in ATP. Although it uses a trajectory region constraining strategy to reduce the location space for perturbation, we omit this strategy and instead use reachability constraint in NGram to illustrate the effect.
  • NGram (Cunningham et al., 2021b): We consider a strong reachability constraint, i.e., from each location, the next location can only be within a distance of $0.75\cdot\hat{a}$, where $\hat{a}$ is the distance between the start and end longitude of the location space. This constraint significantly reduces the location space for perturbation, especially near the boundary.
  • TraCS: We use the same configurations as in the continuous space experiments. For each perturbed location from TraCS, we round it to the nearest discrete location.
The privacy parameter $\varepsilon$ in ATP is for an entire trajectory and can be distributed to each location. For a fair comparison, we use $\varepsilon$ as the privacy parameter for each location. NGram with reachability constraint also uses the same $\varepsilon$ for each location. We omit LDPTrace (Du et al., 2023) as it focuses on trajectory synthesis for a location space rather than perturbation for a given trajectory. The comparison is conducted on the following datasets:
  • Synthetic dataset: We discretize $[0,1)\times[0,1)$ into $m$ discrete points and treat the $m$ points as the location space.
  • NYC and CHI: These real-world datasets are the same as in ATP, extracted from the Foursquare dataset (Yang et al., 2015) and Gowalla dataset (Cho et al., 2011), respectively. [^3] NYC is in $[-74.26,-73.70]\times[40.56,40.97]$ and CHI in $[-87.9,-87.5]\times[41.6,42.0]$. NYC contains $3,478$ discrete locations; CHI contains $1,000$ discrete locations. We use the first $100$ trajectories for evaluation.
For the synthetic dataset, we generate $100$ random trajectories with $100$ locations each to compute the average $AE$. For the NYC and CHI datasets, we perform the perturbation $5$ times for each trajectory and compute the average $AE$.
(a) NYC dataset
(b) CHI dataset
Figure 10. Comparison on real-world datasets.
[…]
Comparison on real-world datasets.
Table 2. Time cost comparison.
ATP NGram TraCS-D TraCS-C
Total (ms) $145.7$ $100.9$ $0.06$ $0.05$
Perturbation $125.8$ $92.8$ $0.018$ $0.003$
4.2.2. Error Comparison on Synthetic Dataset
Figure 9(b) presents the error comparison on the synthetic dataset. Figure 9(a) shows results for $m=10\times 10$, and Figure 9(b) shows results for $m=50\times 50$. We observe that TraCS-C consistently outperforms ATP and NGram across all the $\varepsilon$ values, and TraCS-D also outperforms ATP and NGram when $\varepsilon\gtrsim 4$. Moreover, the performance advantage of TraCS becomes more significant as $m$ increases. This is because finer-grained discretization allows TraCS’s rounding to align with the private locations more accurately. Statistically, the mean $AE$ of TraCS-C is $65.2\%$ of ATP and $77.4\%$ of NGram in Figure 9(a), and $61.5\%$ of ATP and $73.2\%$ of NGram in Figure 9(b). The mean $AE$ of TraCS-D is $75.5\%$ of ATP and $89.6\%$ of NGram in Figure 9(a), and $89.0\%$ of ATP and $74.7\%$ of NGram in Figure 9(b). These results indicate the advantage of TraCS for synthetic trajectories in discrete spaces.
4.2.3. Error Comparison on Real-world Dataset
Figure 10(b) presents the error comparison on the NYC and CHI datasets. ATP and NGram perform better than in the synthetic dataset, especially when $\varepsilon$ is small. However, TraCS still outperforms them as $\varepsilon$ increases.
Reasons for the results. (i) ATP performs exceptionally well when $\varepsilon$ is small, particularly on the NYC dataset. This is because this dataset contains many “condensed” locations, i.e., no other locations near the trajectory. In such cases, if the bi-direction perturbation of ATP is accurate, the perturbed location is error-free. (ii) The $AE$ of ATP does not decrease significantly as $\varepsilon$ increases, which is consistent with their paper. NGram shows a similar trend, especially on the NYC dataset. This is due to the low performance of the Exponential mechanism for large location spaces, particularly when the locations are concentrated in a small area. In such cases, the score functions of the mechanism struggle to differentiate the perturbation quality among the locations. (iii) TraCS exhibits larger $AE$ than ATP and NGram when $\varepsilon$ is small because it focuses on the rectangular area formed by the longitude and latitude, instead of a set of discrete locations. Therefore, the perturbed locations may be far from the city center when $\varepsilon$ is small.
4.2.4. Time Cost Comparison
Table 2 presents the time cost comparison on the NYC dataset. The time cost is measured in milliseconds and averaged per location. It shows that TraCS-D and TraCS-C are significantly faster than ATP and NGram. This efficiency is primarily due to the piecewise-based mechanisms in TraCS, which require negligible time for perturbation compared to the Exponential mechanism in ATP and NGram.
ATP has the highest time cost because it processes two copied trajectories. NGram has a lower time cost than ATP but is still significantly higher than TraCS. These high time costs are primarily due to the Exponential mechanism, which requires multiple evaluations of score functions and sampling from complex distributions (detailed in Appendix B.1). The perturbation procedure accounts for $86.3\%$ and $92.0\%$ of the total time cost in ATP and NGram, respectively.
The total time cost of TraCS is less than $0.5\%$ of ATP and NGram. The perturbation procedure of TraCS-D and TraCS-C only takes $30\%$ and $6\%$ time, respectively. TraCS-D has a higher time cost than TraCS-C because it needs to calculate the distance space $\mathcal{D}_{r(\varphi)}$ for each location in the trajectory, while TraCS-C can use the same $\mathcal{D}_{a}$ and $\mathcal{D}_{b}$ for all locations in the trajectory.
5. Related Work
Trajectory Collection under LDP
Based on the privacy notion, trajectory collection methods under LDP can be categorized into two groups: pure LDP and relaxed LDP. Compared with relaxed LDP, pure LDP imposes a stronger constraint for indistinguishability.
Existing trajectory collection methods under pure LDP include NGram (Cunningham et al., 2021b), LDPTrace (Du et al., 2023), and ATP (Zhang et al., 2023). Among them, LDPTrace focuses on trajectory synthesis within a given location space, while NGram, ATP, and this paper focus on trajectory perturbation.  NGram attempts to ensure pure LDP on the “trajectory space”, which is stronger than the location space. However, this is proved extremely challenging as the trajectory space increases exponentially with the number of possible locations. ATP and this paper focus on ensuring pure LDP in the location space. Unlike ATP and other existing methods, this paper focuses on continuous location spaces, which are stronger than any fine-grained discrete location space.
As a special case of trajectory collection, L-SRR (Wang et al., 2022) focuses on single-location publishing. It discretizes the location space into different-sized cells, with smaller cells near the private location and larger cells farther away. Then, SRR (a variant of $k$-RR) is applied to the cells. Similar to other discretization methods, the number of cells affects its privacy and utility. TraCS provides solutions to this problem in continuous spaces, achieving better privacy and utility, particularly when the number of cells is large.
A widely used relaxed LDP notion for trajectory collection is Geo-indistinguishability (Andrés et al., 2013; Mendes et al., 2020), where the indistinguishability between two locations is a function of their distance, i.e. privacy decreases with distance. Input-discriminative LDP (Gu et al., 2020) goes one step further, assigning each location its privacy level. Threshold-integrated LDP (Yang et al., 2024) defines a threshold to constrain the private region near a location. This is similar to NGram’s reachability constraint but with a formal privacy definition.  More relaxed LDP notions for trajectory collection can be found in (Miranda-Pascual et al., 2023). Although these relaxed LDP notions generally provide better trajectory utility than pure LDP, they sacrifice the strong privacy guarantee of pure LDP, and often need hyperparameters to define the privacy level.
Trajectory Collection under DP
As a weaker privacy notion than LDP, DP (Dwork, 2006; Dwork and Roth, 2014) requires only that any two neighboring locations in the location space be indistinguishable, rather than any two arbitrary locations.
The usage of direction information in trajectory collection dates back to an early work SDD (Jiang et al., 2013), which attempts to add unbounded Laplace noise to the direction information. It provides a DP guarantee for unbounded continuous location spaces, but results in poor utility. The notion of “n-gram” dates back to (Chen et al., 2012), where it abstracts a trajectory pattern as a gram.
External knowledge is also widely used in DP-based trajectory collection methods. NGram’s prior work (Cunningham et al., 2021a) provides a DP-based solution for trajectory synthesis by exploiting the public knowledge of the road network. Focusing on particular privacy-concerned locations, or sensitive zones (Wang and Kankanhalli, 2020), can also improve the trajectory utility while maintaining partial privacy guarantees.
6. Conclusions
This paper presents TraCS-D and TraCS-C, providing local differential privacy guarantees for trajectory collection in continuous location spaces. TraCS-D uses a novel direction-distance perturbation procedure, while TraCS-C perturbs the Cartesian coordinates of each location. These two methods can also be applied to discrete location spaces by rounding the perturbed location to the nearest discrete point. Their trajectory utility is theoretically analyzed and empirically evaluated. Evaluation results on discrete location spaces show that TraCS outperforms state-of-the-art methods in trajectory utility, particularly when the privacy parameter is large.
Beyond trajectory collection, the proposed perturbation mechanisms can also be applied to other scenarios with continuous and bounded domains, providing a direction for future work.
Appendix A Proofs
A.1. Proof of Theorem 3.4
Proof.
$\forall\tau_{1},\tau_{2},\tau^{*}\in\mathcal{S}:\frac{pdf[\mathcal{M}(\tau_{1})=\tau^{*}]}{pdf[\mathcal{M}(\tau_{2})=\tau^{*}]}\leq\exp(\varepsilon).$
In TraCS-D, each location $\tau$ is represented in a direction-distance space with unique coordinates $(\varphi,r(\varphi))$. The key insight of the proof is that the perturbation of $\varphi$ and $r(\varphi)$ satisfies $\varepsilon_{d}$-LDP and $(\varepsilon-\varepsilon_{d})$-LDP, respectively. Then TraCS-D ensures $\varepsilon$-LDP for the location space $\mathcal{S}=\mathcal{D}_{\varphi}\otimes\mathcal{D}_{r(\varphi)}$.
The definitions of mechanisms $\mathcal{M}_{\circ}$ and $\mathcal{M}_{-}$ imply that they satisfy LDP. Specifically, for $\mathcal{M}_{\circ}(\varphi;\varepsilon_{d})$,
$\forall\varphi_{1},\varphi_{2},\varphi^{*}\in[0,2\pi):\frac{pdf[\mathcal{M}_{\circ}(\varphi_{1})=\varphi^{*}]}{pdf[\mathcal{M}_{\circ}(\varphi_{2})=\varphi^{*}]}\leq\exp(\varepsilon_{d}),$
is guaranteed by Definition 3.1 of $\mathcal{M}_{\circ}$. Similarly, $\mathcal{M}_{-}(\overline{r}(\varphi);\varepsilon-\varepsilon_{d})$ in definition 3.3 ensures $(\varepsilon-\varepsilon_{d})$-LDP for $\overline{r}(\varphi)$, i.e.
$\forall\overline{r}_{1},\overline{r}_{2},\overline{r}^{*}\in[0,1):\frac{pdf[\mathcal{M}_{-}(\overline{r}_{1})=\overline{r}^{*}]}{pdf[\mathcal{M}_{-}(\overline{r}_{2})=\overline{r}^{*}]}\leq\exp(\varepsilon-\varepsilon_{d}).$
$\forall r_{1},r_{2},r^{*}\in\mathcal{D}_{r(\varphi^{\prime})}:\frac{pdf[r_{1}=r^{*}]}{pdf[r_{2}=r^{*}]}\leq\exp(\varepsilon-\varepsilon_{d}),$
because the linear mapping $g$ is a deterministic function without randomness. Therefore, $\mathcal{M}_{-}$ ensures $(\varepsilon-\varepsilon_{d})$-LDP for any $\mathcal{D}_{r(\varphi^{\prime})}$.
Combining these results with Sequential Composition Theorem 2.2, we can conclude that $\mathcal{M}=(\mathcal{M}_{\circ},\mathcal{M}_{-})$ ensures $\varepsilon$-LDP for each location $\tau=(\varphi,r(\varphi))\in\mathcal{S}$. Consequently, the entire trajectory with $n$ locations satisfies $n\varepsilon$-LDP. ∎
A.2. Proof of Theorem 3.5
Proof.
[…]
The worst-case MSE is achieved when $\overline{r}=0$ or $\overline{r}=1$, i.e. at the endpoints (similar to the variance calculation in (Wang et al., 2019)). In this case,
[…]
with $p_{\varepsilon}$ and $C$ defined in Definition 3.3. We omit the calculation for $\overline{r}=1$ as it is symmetric to $\overline{r}=0$. Plugging in the values of $p_{\varepsilon}$ and $C$, we can simplify the MSE using big $\mathcal{O}$ notation:
A.3. Proof of Theorem 3.6
Proof.
In Proof A.1, we have already proved that $\mathcal{M}_{-}(\overline{a},\varepsilon/2)$ satisfies $\varepsilon/2$-LDP for the normalized distance $\overline{a}$. Then, the linear mapping from $\overline{a}\in[0,1)$ to $a\in\mathcal{D}_{a}$ is a post-processing step that does not affect the LDP property. This means $\mathcal{M}_{-}$ ensures $\varepsilon/2$-LDP for the distance space $\mathcal{D}_{a}$. In the same way, it also ensures $\varepsilon/2$-LDP for the distance space $\mathcal{D}_{b}$. Therefore, $\mathcal{M}=(\mathcal{M}_{-},\mathcal{M}_{-})$ satisfies $\varepsilon$-LDP for each location $\tau=(a,b)\in\mathcal{S}$ by Sequential Composition Theorem 2.2. Consequently, the entire trajectory with $n$ locations satisfies $n\varepsilon$-LDP. ∎
Appendix B Complementary Materials
B.1. Weakness of the Exponential Mechanism
This section discusses the Exponential mechanism and its weaknesses in the context of trajectory collection.
Definition B.1 (Exponential Mechanism (Nissim et al., 2012)).
Given a scoring function $d:\mathcal{X}\times\mathcal{Y}\to\mathbb{R}$, a privacy parameter $\varepsilon>0$, and a set of possible outputs $\mathcal{Y}$, the Exponential mechanism $\mathcal{M}_{\exp}:\mathcal{X}\to\mathcal{Y}$ is defined by:
$\Pr[\mathcal{M}_{\exp}(x)=y]=\frac{\exp\left(\frac{\varepsilon d(x,y)}{2\Delta d}\right)}{\sum_{y^{\prime}\in\mathcal{Y}}\exp\left(\frac{\varepsilon d(x,y^{\prime})}{2\Delta d}\right)},$
where $\Delta d=\max_{y,y^{\prime}\in\mathcal{Y}}|d(x,y)-d(x,y^{\prime})|$ is the sensitivity of the scoring function $d$.
In the context of trajectory collection, the most used scoring function is the negative Euclidean distance $d(x,y)\coloneq-||x-y||_{2}$. Therefore, locations with a smaller Euclidean distance to the private location $x$ are more likely to be sampled.
The biggest weakness of the Exponential mechanism is the complexity in calculating the score function and sampling from the probability distribution. It requires computing the Euclidean distance between every pair of locations in the location space, which is computationally expensive for large-scale spaces.  In ATP and NGram, the location space for perturbation changes at each time step, meaning that the score functions need to be recomputed at each perturbation, resulting in $\mathcal{O}(m)$ time complexity for each perturbation, where $m$ is the size of the location space.
The complexity of sampling from the Exponential mechanism is also $\mathcal{O}(m)$, as it needs to sample from an $m$-piece cumulative distribution function (CDF) at each perturbation.
B.2. Dominant Sector Comparison Between the Strawman Approach and TraCS-D
For direction perturbation, a small-size and high-probability dominant sector is desirable for more accurate perturbation. However, these two properties are generally conflicting with each other: a smaller dominant sector typically has a lower probability.  We will show that TraCS-D achieves a better trade-off between the size and probability of the dominant sector compared to the strawman approach with any $k$ value. Figure 11(b) shows two exemplary quantitative comparisons between them.
Small $\varepsilon$ region. In this case, the dominant sector of TraCS-D is larger than the strawman approach, but the probability of the dominant sector is significantly higher, especially when $k$ is large.
We can observe the effect of $k$ on the strawman approach. Large $k$ values reduce the size difference of the dominant sector size compared with TraCS-D, but they also dramatically reduce the probability of the dominant sector in the small $\varepsilon$ region. Conversely, small $k$ values increase the probability of the dominant sector in the small $\varepsilon$ region, but they also lead to a larger dominant sector in the large $\varepsilon$ region.
(a) $k=6$
(b) $k=12$
Figure 11. Comparison of the size and probability of the dominant sector between TraCS-D and the strawman approach.
[…]
Comparison of the size and probability of the dominant sector between TraCS-D and the strawman approach.
B.3. Detailed Expression of $|\mathcal{D}_{r(\varphi)}|$
Then, $|\mathcal{D}_{r(\varphi)}|$ has the following four cases:
depending on the direction $\varphi$ from reference location $\tau_{i}=(a_{i},b_{i})$.
B.4. Redesigned SW
We redesign the SW mechanism to achieve $\varepsilon$-LDP for the circular space $[0,2\pi)$ and the linear space $[0,1)$. The key idea is to transform the original sampling distribution to these spaces while preserving the LDP constraint.  Similar to Definition 3.1, the redesigned SW mechanism for direction perturbation is defined as follows.
Definition B.2 (Redesigned SW for direction perturbation).
Given a private direction $\varphi$ and a privacy parameter $\varepsilon$, redesigned SW for direction perturbation is a mechanism $\mathcal{M}_{\circ}:[0,2\pi)\to[0,2\pi)$ defined by:
@@ B.4. Redesigned SW
where $p_{\varepsilon}=\frac{1}{2\pi\varepsilon}(\exp(\varepsilon)-1)$ is the sampling probability, and $[l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})$ is the sampling interval that
@@
Similar to Definition 3.3, the redesigned SW mechanism for distance perturbation is defined as follows.
Definition B.3 (Redesigned SW for distance perturbation).
Given a private distance $\overline{r}(\varphi)$ and a privacy parameter $\varepsilon$, redesigned SW for distance perturbation is a mechanism $\mathcal{M}_{-}:[0,1)\to[0,1)$ that
@@
Compared with Definition 3.1 and Definition 3.3, the redesigned SW mechanisms have a larger $p_{\varepsilon}$ and a narrower $[l,r)$ or $[u,v]$ interval for sampling. We can treat the redesigned SW mechanisms as a more aggressive perturbation mechanism: it tries to sample from a narrow dominant sector with a higher probability ($p_{\varepsilon}$), which also leads to a higher probability ($p_{\varepsilon}/\exp(\varepsilon)$) of non-dominant sectors due to the LDP constraint.

### v2

@@ C.4.1 Other Piecewise-based Mechanisms
In Definition 5 and Definition 6, we use specific instantiations for the parameters $p$ and $[l,r)$ in the piecewise-based mechanism. Besides these instantiations, TraCS can be integrated with any other piecewise-based mechanisms [31, 21, 24, 23, 38] by redesigning them. For example, SW [21] is a famous mechanism proposed for distribution estimation; it is defined on $0,1)\to[-b,1+b)$, with $b$ as a parameter. It uses different $p$ and $[l,r)$, and employs a strategy of maximizing the mutual information between the perturbed and sensitive data to determine $b$. We can redesign SW for the direction perturbation and distance perturbation in TraCS-D and TraCS-C. Appendix C.5 provides the detailed redesign of SW for them and shows intuitive comparisons with Definition 5 and Definition 6. The experimental section also provides an comparison between the redesigned SW and TraCS-D.
C.4.2 Other Shapes of Location Space
@@
TraCS-D for circular area. When the location space is circular with radius $R$, TraCS-D has a more straightforward design. The main insight is that such a location space $\mathcal{S}_{R}$ has a natural representation in polar coordinates, i.e. $\mathcal{S}_{R}=\{(\varphi,r):\varphi\in[0,2\pi),r\in[0,R)\}$. Any location $\tau_{i}\in\mathcal{S}_{R}$ corresponds to a pair of coordinates $(\varphi_{i},r_{i})$, and this representation is unique and reversible. In this case, the direction space $\mathcal{D}_{\varphi}=[0,2\pi)$ and the distance space $\mathcal{D}_{r}=[0,R)$ are independent, so we can apply $\mathcal{M}_{\circ}$ to $\varphi_{i}$ and $\mathcal{M}_{-}$ to $r_{i}/R$. The perturbed location $\tau_{i}^{\prime}$ is then $(\varphi_{i}^{\prime},r_{i}^{\prime})=(\mathcal{M}_{\circ}(\varphi_{i}),\mathcal{M}_{-}(r_{i}/R)\cdot R)$, still within polar coordinates.
C.5 Redesigned SW (Appendix C.4.1)
We redesign the SW mechanism to achieve $\varepsilon$-LDP for the circular space $0,2\pi)$ and the linear space $[0,1)$. The key idea is to transform the original sampling distribution to these spaces while preserving the LDP constraint. Similar to Definition 5, the redesigned SW mechanism for direction perturbation is defined as follows.
Definition 7 (Redesigned SW for direction perturbation).
Given a sensitive direction $\varphi$ and a privacy parameter $\varepsilon$, redesigned SW for direction perturbation is a mechanism $\mathcal{M}_{\circ}:[0,2\pi)\to[0,2\pi)$ defined by:
@@
where $p_{\varepsilon}=\frac{1}{2\pi\varepsilon}(\exp(\varepsilon)-1)$ is the sampling probability, and $l_{\varphi,\varepsilon},r_{\varphi,\varepsilon})$ is the sampling interval that
@@
Similar to Definition 6, the redesigned SW mechanism for distance perturbation is defined as follows.
Definition 8 (Redesigned SW for distance perturbation).
Given a sensitive distance $\overline{r}(\varphi)$ and a privacy parameter $\varepsilon$, redesigned SW for distance perturbation is a mechanism $\mathcal{M}_{-}:[0,1)\to[0,1)$ that
@@
Compared with Definition 5 and Definition 6, the redesigned SW mechanisms have a larger $p_{\varepsilon}$ and a narrower $[l,r)$ or $[u,v]$ interval for sampling. We can treat the redesigned SW mechanisms as a more aggressive perturbation mechanism: it tries to sample from a narrow dominant sector with a higher probability ($p_{\varepsilon}$), which also leads to a higher probability ($p_{\varepsilon}/\exp(\varepsilon)$) of non-dominant sectors due to the LDP constraint.
C.6 Experimental Results for Extensions (Appendix C.4)
C.6.1 Effect of Different Piecewise-based Mechanisms
Figure 14 compares the performance of different piecewise-based mechanisms for TraCS-D. In addition to the mechanisms defined in Definition 5 and Definition 6, we name the redesigned SW for TraCS-D as TraCS-D (R-SW), as detailed in Appendix C.5. We can observe that TraCS-D consistently exhibits smaller errors than TraCS-D (R-SW) across all the $\varepsilon$ values. Statistically, the mean $AE$ of TraCS-D is $52.6\%$ of TraCS-D (R-SW) across all the $\varepsilon$ values. This difference comes from the MSE of their sampling distributions, where Definition 5 and Definition 6 have smaller MSE than R-SW.
C.6.2 TraCS-D for Circular Area
Figure 14 illustrates an example of TraCS-D applied to a circular area $\mathcal{S}=[0,2\pi)\otimes[0,1)$. We set the sensitive location $\tau$ as $(\pi,0.5)$ and the privacy parameter $\varepsilon=5$, then collect $50$ random samples of perturbed locations using TraCS-D. In this case, the dominant area is $0.87\pi,1.13\pi)\otimes[0.32,0.67)$, as determined by the perturbation mechanisms in Definition 5 and Definition 6. We can observe that the perturbed locations are concentrated in the dominant area.
Figure 13: Comparison between different piecewise-based mechanisms.
Figure 14: Fifty samples of TraCS-D at $\tau=(\pi,0.5)$ for circular area.
C.7 Experimental Results for Other Metrics (Section 4.1.1)
A smaller $AE$ generally indicates better preservation of range queries and hotspots [9, 37]. (i) For range query preservation (Formula 17 in [9]), given a threshold $\delta$, a perturbed location is considered correct if it lies within a $\delta$ distance of the corresponding sensitive location. Thus, when $AE\leq\delta$, the perturbed location is expected to satisfy the range query, and a smaller $AE$ implies higher correctness. (ii) For hotspot preservation (Section 6.2.4 in [37]), which measures how many hotspots from a given set remain after perturbation, a smaller $AE$ leads to better hotspot retention after rounding in TraCS.
(a) Range query preservation
(b) Hotspot preservation
Figure 15: Comparison of other trajectory metrics on the CHI dataset. Higher range query preservation and lower error in hotspot preservation indicate better performance.
We further evaluate the performance of TraCS-D and existing methods under varying privacy parameters $\varepsilon$ using these two metrics. Specifically, we use the CHI dataset and adopt the following experimental setups:
  • Range query preservation: We set the threshold $\delta=0.1$, meaning a perturbed location is considered correct if its distance from the sensitive location is less than $0.1$ in longitude-latitude coordinates.
  • Hotspot preservation: The hotspot set is defined as the first $50\%$ of locations in the CHI dataset. We count the number of hotspots before and after perturbation for the first $100$ trajectories and report the difference.
The results are shown in Figure 15. It can be seen that TraCS generally outperforms existing methods across different privacy levels, demonstrating superior effectiveness in preserving both range queries and hotspots. Particularly, TraCS shows significant improvement in hotspot preservation compared to existing methods across all privacy levels.